# Search for a diffuse flux of photons with energies above tens of PeV at the Pierre Auger Observatory

### v2

@@ 6 Results of the photon search
Excluding the burnt sample, the remaining $14,299$ events constitute the search data set. For each event, the $M_{1}$ observable is compared to the parametrized photon candidate cut $M_{1}^{\text{cut}}$ for its respective event category, as shown in Fig. 13. The parameterizations of $M_{1}^{\text{cut}}$, performed with simulated events up to $10^{17.5}$ eV, have been extrapolated to higher energies for comparison with the data. Nevertheless, all events exhibit $M_{1}$ values well above the parametrized $M_{1}^{\text{cut}}$, indicating that no photon candidate events are identified.
@@ 6.1 Upper limits calculation
Figure 13: The observable $M_{1}$ in terms of the photon-equivalent energy for events in the search data set. Each panel corresponds to the quoted categories as listed in Table 4 from left to right, and top to bottom. Solid lines represent the parametrized photon candidate cut for each category.
The accumulated exposure to photon primaries is derived from the continuous, real-time monitoring of the SD-433 array, as described in Section 2. The operational status of each SD station is known every second, allowing stations to be classified as either “on” or “off”. This enables real-time monitoring of the status of each hexagon [^6] Technical issues affecting eight out of the $19$ SD stations during $16$ distinctive periods (lasting between one day and one week) were identified, accounting for less than $1\%$ of the operation time. During these periods, the stations were considered “off”.. Following the 5T5 and 6T5 event selection criteria, a hexagon is considered active if at least five “on” WCDs are surrounding the hottest one. When a 6T5 condition is met, the effective area of a hexagon with a side length $d=433\,\mathrm{m}$ is the one of the unitary hexagon, i.e., $A=\sqrt{3}/2\,\times\,d^{2}$. For the 5T5 condition, the area is reduced by a geometrical factor of $1/3$ to account for the missing nearest-neighbor station. Consequently, an active hexagon contributes to the exposure calculation with a time-dependent area, $A(t)$, reflecting its real-time status.
As discussed in Section 5, a minimum area spanned by UMD modules is required to adequately sample the air-shower muon content. The number of “on” UMD modules at the central station and the first ring around it, as well as the relative position of any missing UMD stations in the first ring, are dynamic. They vary over time due to the absence of deployed modules or technical issues affecting normal operation. These requirements are mapped by a step function, $u(t)$, incorporated into the exposure calculation.
The trigger efficiency for photon-initiated showers, $\tau$, must also be included in the exposure calculation. This efficiency depends on the primary energy and zenith angle, as discussed in Section 3.2, and the instantaneous exposure decreases as the primary energy decreases.
@@
In Fig. 14, we present the upper limits on the integral photon flux obtained in this study (represented by red markers), alongside the limits previously reported by the Auger Collaboration [71, 35], by KASCADE-Grande [21], by EAS-MSU [22], and by Telescope Array [72, 73]. The limits derived in this work are the only ones based on measurements from the Southern Hemisphere in the tens of PeV energy domain. This analysis extends the photon search program of Auger, covering over three decades of cosmic-ray energy. It is the first one in which direct measurements of air-shower muons provided by the UMD are employed.
Diffuse photon fluxes are expected from the interaction of UHE cosmic rays with background radiation fields [74, 8, 7] and with the interstellar Galactic matter [75], as discussed in Section 1. These are indicated by the shaded bands in Fig. 14. The upper limits derived in this study are between two and three orders above these cosmogenic fluxes. However, the expected diffuse gamma-ray flux from proton-proton interactions in the galactic halo [76] is shown to be within the reach of the upper limits obtained in this work. Moreover, predictions from various phenomena beyond the Standard Model may emerge above these cosmogenic fluxes, e.g., those arising from the decay of super-heavy dark matter particles [26]. In the above-mentioned figure, we show the predicted diffuse fluxes in several scenarios: assuming decay channels into hadrons, a mass $M_{\text{X}}=10^{10}\,\mathrm{GeV}$ and a lifetime $\tau_{\text{X}}=3\times 10^{21}\,\mathrm{yr}$ [77]; $M_{\text{X}}=10^{12}\,\mathrm{GeV}$ and $\tau_{\text{X}}=10^{23}\,\mathrm{yr}$ [77]; and assuming decay into leptons, a mass $M_{\text{X}}=10^{10}\,\mathrm{GeV}$ and a lifetime $\tau_{\text{X}}=3\times 10^{21}\,\mathrm{yr}$ [27]. Further sensitivity provided by the next 10 years of data taking will allow us to constrain the parameter space of the discussed models thanks to an expected improvement in the upper limits by a factor larger than $20$.
@@ 6.2 Systematic uncertainties on the upper limits
Proton events are assigned an overestimated energy in the photon-equivalent scale, with an energy bias that decreases from $10\%$ to $5\%$ in the energy range of interest (see Fig. 6, right). Following Eq. 6.3, the effect of this bias on the upper limits ranges from $-10\%$ to $-6\%$ with increasing energy. Since this would imply a decrease in the upper limits, we neglect this effect to maintain a conservative calculation. Similarly, a negative energy bias between $-10\%$ and $-15\%$, decreasing linearly with the logarithm of energy, is found for photon events in the photon-equivalent energy scale (see Fig. 6, right). This bias leads to a relative increase of the upper limits, ranging from $12\%$ to $17\%$ with increasing energy.
@@
The impact of variations in the spectral index assumed in the energy-integrated exposure calculation has been examined. A variation of ±$\pm 0.5$ in the spectral index results in an average change in the exposure of between $0.7\%$ and $1.7\%$ for the quoted threshold energies. Consequently, the upper limits on the integral flux increase by a factor ranging from $0.8\%$ to $2.2\%$ with increasing energy.
@@
The number of photon candidate events can increase if the values of $M_{1}$ decrease. In the following, we discuss how varying the parameters in the $M_{1}$ definition might cause this effect. First, the measured muon density can decrease by up to $5.6\%$ when varying the parameters involved in the muon counting strategy used to convert digital traces to the number of muons in the UMD stations [46]. Second, the UMD simulation assumes a soil density of 2.38 g cm-3, leading to a fixed soil shielding to muons. The standard deviation of the soil density is 0.05 g cm-3 [78], which results in a $2.8\%$ increase in the reference muon density $\rho_{\text{pr}}$ [42]. Lastly, different high-energy hadronic models predict varying numbers of muons in simulations, which could affect $\rho_{\text{pr}}$. The only model surpassing the employed one, EPOS-LHC, in terms of the predicted number of muons, is Sibyll2.3d by about $5\%$ [79]. Combining the three factors, the overall systematic effect translates into an absolute shift of $-0.06$ across all $M_{1}$ values. Despite this shift, no event falls below the parametrized photon candidate cuts.
@@
Following the conservative approach mentioned before, the upper limits above each threshold energy can be increased by the corresponding factors arising from the bias in the photon-equivalent energy and the spectral index employed in the exposure calculation. Thus, the final upper limits reported in this work, including a shift towards the worst-case scenario motivated by the considered systematic effects, are $13.8$, $13.5$, $13.3$ and $13.6$ km-2 sr-1 yr-1 for threshold energies of $50$, $80$, $120$ and 200 PeV. A summary of the upper limits before and after accounting for the systematic effects is presented in Table 5. These limits translate into constraints on the photon fraction in the measured cosmic-ray flux [80]. Above the specified energy thresholds, the upper limits on the photon fraction are $0.056\%$, $0.14\%$, $0.35\%$ and $0.97\%$ at a confidence level $95\%$.
@@
Table 5: The upper limits on the integral photon flux above each quoted energy threshold before (third column) and after (fourth column) accounting for the discussed systematic effects. The exposure above each energy threshold is obtained from Eq. 6.1.
@@ Appendix A Lateral distribution function for photon events
where the reference zenith angle, 25 , is selected as the median of the $E_{\text{MC}}^{-2}$-weighted zenith distribution. The three free parameters of $\tilde{f}(\theta)$ are estimated by examining the approximately linear relationship between the $S_{250}$ and $E_{\text{MC}}$, as shown in Fig. A.1, left. Consequently, the ratio $S(250)/E_{\text{MC}}$ is mainly determined by the angular dependence, as depicted in Fig. A.1, right. The sigmoid function $\tilde{f}(\theta)$ is fitted to the average ratio, resulting in parameters $f_{0}=(4.65\pm 0.03)\,\mathrm{VEM/10^{16}\,\mathrm{eV}}$, $b=(0.194\pm 0.002)$ and $c=(0.156\pm 0.002)$. This model captures the mean behavior within a $2\%$ margin.
Figure A.1: Left: The shower size, $S(250)$, as a function of the simulated primary energy, $E_{\text{MC}}$, for photon events. The solid lines represent linear fits in zenith-angle bins. Right: The average ratio between the shower size and the simulated energy versus the zenith angle. The solid line corresponds to the fit $\tilde{f}(\theta)$ from Eq. A.3 via a $\chi^{2}$ minimization.
The linear coefficients, $\beta_{0}$ and $\beta_{1}$, depend quadratically on $\sec\theta$, as illustrated in Fig. A.2, left. The corresponding six parameters are estimated using an unbinned maximum likelihood method, assuming a Gaussian probability density function for $\beta$. The resulting values are given in Eq. A.4. The performance of the parametrization is assessed by comparing the predicted slope from Eq. A.2, $\beta_{\text{pred}}$, with the reconstructed slope for each photon-initiated event, $\beta$. The parametrization shows a negligible bias (within ±$\pm 2\%$) and a resolution better than $19\%$ as illustrated in Fig. A.2, right.
Figure A.2: Left: The evolution of $\beta_{0}$ and $\beta_{1}$ parameters of Eq. A.2 with $\sec\theta$. The second-order polynomials in Eq. A.4 are superimposed. Right: The bias and resolution of the LDF slope predicted by Eq. A.2, $\beta_{\text{pred}}$.
$\beta_{0}$ $=(0.756\pm 0.011)+(4.05\pm 0.01)\times\sec\theta-(2.33\pm 0.01)\times\sec^{2}\theta$ (A.4)
$\beta_{1}$ $=(0.704\pm 0.006)-(1.37\pm 0.01)\times\sec\theta+(0.758\pm 0.004)\times\sec^{2}\theta$
@@ Appendix B Parametrization of the background contamination and photon candidate cuts
The average background contamination at $50\%$ signal efficiency and the photon candidate cut were parametrized in terms of the photon-equivalent energy scale. As shown by the dashed line in both panels of Fig. 11, the employed analytical models are:
@@
The free parameters of both equations are given in Table A.1 for each event category.
Event category $b_{0}$ $b_{1}$ $c_{0}$ $c_{1}$ $c_{2}$
@@
Table A.1: The parameters of Eqs. B.1 and B.2 modeling the background contamination at $50\%$ signal efficiency and the photon candidate cut for each event category.

### v3

@@ 6 Results of the photon search
Excluding the burnt sample, the remaining $14,299$ events constitute the search data set. For each event, the $M_{1}$ observable is compared to the parametrized photon candidate cut $M_{1}^{\text{cut}}$ for its respective event category, as shown in Fig.˜13. The parameterizations of $M_{1}^{\text{cut}}$, performed with simulated events up to $10^{17.5}$ eV, have been extrapolated to higher energies for comparison with the data. Nevertheless, all events exhibit $M_{1}$ values well above the parametrized $M_{1}^{\text{cut}}$, indicating that no photon candidate events are identified.
@@ 6.1 Upper limits calculation
Figure 13: The observable $M_{1}$ in terms of the photon-equivalent energy for events in the search data set. Each panel corresponds to the quoted categories as listed in Table˜4 from left to right, and top to bottom. Solid lines represent the parametrized photon candidate cut for each category.
The accumulated exposure to photon primaries is derived from the continuous, real-time monitoring of the SD-433 array, as described in Section˜2. The operational status of each SD station is known every second, allowing stations to be classified as either “on” or “off”. This enables real-time monitoring of the status of each hexagon [^6] Technical issues affecting eight out of the $19$ SD stations during $16$ distinctive periods (lasting between one day and one week) were identified, accounting for less than $1\%$ of the operation time. During these periods, the stations were considered “off”.. Following the 5T5 and 6T5 event selection criteria, a hexagon is considered active if at least five “on” WCDs are surrounding the hottest one. When a 6T5 condition is met, the effective area of a hexagon with a side length $d=433\,\mathrm{m}$ is the one of the unitary hexagon, i.e., $A=\sqrt{3}/2\,\times\,d^{2}$. For the 5T5 condition, the area is reduced by a geometrical factor of $1/3$ to account for the missing nearest-neighbor station. Consequently, an active hexagon contributes to the exposure calculation with a time-dependent area, $A(t)$, reflecting its real-time status.
As discussed in Section˜5, a minimum area spanned by UMD modules is required to adequately sample the air-shower muon content. The number of “on” UMD modules at the central station and the first ring around it, as well as the relative position of any missing UMD stations in the first ring, are dynamic. They vary over time due to the absence of deployed modules or technical issues affecting normal operation. These requirements are mapped by a step function, $u(t)$, incorporated into the exposure calculation.
The trigger efficiency for photon-initiated showers, $\tau$, must also be included in the exposure calculation. This efficiency depends on the primary energy and zenith angle, as discussed in Section˜3.2, and the instantaneous exposure decreases as the primary energy decreases.
@@
In Fig.˜14, we present the upper limits on the integral photon flux obtained in this study (represented by red markers), alongside the limits previously reported by the Auger Collaboration [71, 35], by KASCADE-Grande [21], by EAS-MSU [22], and by Telescope Array [72, 73]. The limits derived in this work are the only ones based on measurements from the Southern Hemisphere in the tens of PeV energy domain. This analysis extends the photon search program of Auger, covering over three decades of cosmic-ray energy. It is the first one in which direct measurements of air-shower muons provided by the UMD are employed.
Diffuse photon fluxes are expected from the interaction of UHE cosmic rays with background radiation fields [74, 8, 7] and with the interstellar Galactic matter [75], as discussed in Section˜1. These are indicated by the shaded bands in Fig.˜14. The upper limits derived in this study are between two and three orders above these cosmogenic fluxes. However, the expected diffuse gamma-ray flux from proton-proton interactions in the galactic halo [76] is shown to be within the reach of the upper limits obtained in this work. Moreover, predictions from various phenomena beyond the Standard Model may emerge above these cosmogenic fluxes, e.g., those arising from the decay of super-heavy dark matter particles [26]. In the above-mentioned figure, we show the predicted diffuse fluxes in several scenarios: assuming decay channels into hadrons, a mass $M_{\text{X}}=10^{10}\,\mathrm{GeV}$ and a lifetime $\tau_{\text{X}}=3\times 10^{21}\,\mathrm{yr}$ [77]; $M_{\text{X}}=10^{12}\,\mathrm{GeV}$ and $\tau_{\text{X}}=10^{23}\,\mathrm{yr}$ [77]; and assuming decay into leptons, a mass $M_{\text{X}}=10^{10}\,\mathrm{GeV}$ and a lifetime $\tau_{\text{X}}=3\times 10^{21}\,\mathrm{yr}$ [27]. Further sensitivity provided by the next 10 years of data taking will allow us to constrain the parameter space of the discussed models thanks to an expected improvement in the upper limits by a factor larger than $20$.
@@ 6.2 Systematic uncertainties on the upper limits
Proton events are assigned an overestimated energy in the photon-equivalent scale, with an energy bias that decreases from $10\%$ to $5\%$ in the energy range of interest (see Fig.˜6, right). Following Eq.˜6.3, the effect of this bias on the upper limits ranges from $10\%$ to $6\%$ with increasing energy. Since this would imply a decrease in the upper limits, we neglect this effect to maintain a conservative calculation. Similarly, a negative energy bias between $10\%$ and $15\%$, decreasing linearly with the logarithm of energy, is found for photon events in the photon-equivalent energy scale (see Fig.˜6, right). This bias leads to a relative increase of the upper limits, ranging from $12\%$ to $17\%$ with increasing energy.
@@
The impact of variations in the spectral index assumed in the energy-integrated exposure calculation has been examined. A variation of $\pm 0.5$ in the spectral index results in an average change in the exposure of between $0.7\%$ and $1.7\%$ for the quoted threshold energies. Consequently, the upper limits on the integral flux increase by a factor ranging from $0.8\%$ to $2.2\%$ with increasing energy.
@@
The number of photon candidate events can increase if the values of $M_{1}$ decrease. In the following, we discuss how varying the parameters in the $M_{1}$ definition might cause this effect. First, the measured muon density can decrease by up to $5.6\%$ when varying the parameters involved in the muon counting strategy used to convert digital traces to the number of muons in the UMD stations [46]. Second, the UMD simulation assumes a soil density of 2.38 g cm-3, leading to a fixed soil shielding to muons. The standard deviation of the soil density is 0.05 g cm-3 [78], which results in a $2.8\%$ increase in the reference muon density $\rho_{\text{pr}}$ [42]. Lastly, different high-energy hadronic models predict varying numbers of muons in simulations, which could affect $\rho_{\text{pr}}$. The only model surpassing the employed one, EPOS-LHC, in terms of the predicted number of muons, is Sibyll2.3d by about $5\%$ [79]. Combining the three factors, the overall systematic effect translates into an absolute shift of 0.06-0.06 across all $M_{1}$ values. Despite this shift, no event falls below the parametrized photon candidate cuts.
@@
Following the conservative approach mentioned before, the upper limits above each threshold energy can be increased by the corresponding factors arising from the bias in the photon-equivalent energy and the spectral index employed in the exposure calculation. Thus, the final upper limits reported in this work, including a shift towards the worst-case scenario motivated by the considered systematic effects, are $13.8$, $13.5$, $13.3$ and $13.6$ km-2 sr-1 yr-1 for threshold energies of $50$, $80$, $120$ and 200 PeV. A summary of the upper limits before and after accounting for the systematic effects is presented in Table˜5. These limits translate into constraints on the photon fraction in the measured cosmic-ray flux [80]. Above the specified energy thresholds, the upper limits on the photon fraction are $0.056\%$, $0.14\%$, $0.35\%$ and $0.97\%$ at a confidence level $95\%$.
@@
Table 5: The upper limits on the integral photon flux above each quoted energy threshold before (third column) and after (fourth column) accounting for the discussed systematic effects. The exposure above each energy threshold is obtained from Eq.˜6.1.
@@ Appendix A Lateral distribution function for photon events
where the reference zenith angle, 25 , is selected as the median of the $E_{\text{MC}}^{-2}$-weighted zenith distribution. The three free parameters of $\tilde{f}(\theta)$ are estimated by examining the approximately linear relationship between the $S_{250}$ and $E_{\text{MC}}$, as shown in Fig.˜A.1, left. Consequently, the ratio $S(250)/E_{\text{MC}}$ is mainly determined by the angular dependence, as depicted in Fig.˜A.1, right. The sigmoid function $\tilde{f}(\theta)$ is fitted to the average ratio, resulting in parameters $f_{0}=(4.65\pm 0.03)\,\mathrm{VEM/10^{16}\,\mathrm{eV}}$, $b=(0.194\pm 0.002)$ and $c=(0.156\pm 0.002)$. This model captures the mean behavior within a $2\%$ margin.
Figure A.1: Left: The shower size, $S(250)$, as a function of the simulated primary energy, $E_{\text{MC}}$, for photon events. The solid lines represent linear fits in zenith-angle bins. Right: The average ratio between the shower size and the simulated energy versus the zenith angle. The solid line corresponds to the fit $\tilde{f}(\theta)$ from Eq.˜A.3 via a $\chi^{2}$ minimization.
The linear coefficients, $\beta_{0}$ and $\beta_{1}$, depend quadratically on $\sec\theta$, as illustrated in Fig.˜A.2, left. The corresponding six parameters are estimated using an unbinned maximum likelihood method, assuming a Gaussian probability density function for $\beta$. The resulting values are given in Eq.˜A.4. The performance of the parametrization is assessed by comparing the predicted slope from Eq.˜A.2, $\beta_{\text{pred}}$, with the reconstructed slope for each photon-initiated event, $\beta$. The parametrization shows a negligible bias (within $\pm 2\%$) and a resolution better than $19\%$ as illustrated in Fig.˜A.2, right.
Figure A.2: Left: The evolution of $\beta_{0}$ and $\beta_{1}$ parameters of Eq.˜A.2 with $\sec\theta$. The second-order polynomials in Eq.˜A.4 are superimposed. Right: The bias and resolution of the LDF slope predicted by Eq.˜A.2, $\beta_{\text{pred}}$.
[…]
@@ Appendix B Parametrization of the background contamination and photon candidate cuts
The average background contamination at $50\%$ signal efficiency and the photon candidate cut were parametrized in terms of the photon-equivalent energy scale. As shown by the dashed line in both panels of Fig.˜11, the employed analytical models are:
@@
The free parameters of both equations are given in Table˜A.1 for each event category.
Event
category
$b_{0}$ $b_{1}$ $c_{0}$ $c_{1}$ $c_{2}$
@@
Table A.1: The parameters of Eqs.˜B.1 and B.2 modeling the background contamination at $50\%$ signal efficiency and the photon candidate cut for each event category.